OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points)

OpenStudy (homeworkhelper099):

your trying to solve for X.

OpenStudy (anonymous):

yes but i have to explain how i can show all the solutions for that equation and i get so lost

OpenStudy (homeworkhelper099):

oh i see. let me think and i'll get back with you..

OpenStudy (dtan5457):

If you simplify this out to a function you get y=9x+2. To get solution pairs (which are infinite), you just put any x value, to get a corresponding y value. "Part B: Determine three different solutions for this equation. (4 points)" I'll give you one solution f(1)= 9(1)+2=11 (1,11) Inconsistent lines have the same slope, but different y intercepts.

OpenStudy (anonymous):

ok so thats like the answer for part a:

OpenStudy (anonymous):

im so lost .-.

OpenStudy (dtan5457):

Alright so for part B Part B: Determine three different solutions for this equation. (4 points) I'll give you one solution If I wanted x=1 9(1)+2=11 (1,11)

OpenStudy (dtan5457):

Create 2 more

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so 9(2)+2=20 and 9(3)+2=29

OpenStudy (dtan5457):

yes

OpenStudy (dtan5457):

now for part c "Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent." y=9x+2 is in the form of y=mx+b m=slope b=y intercept if you want them to be inconsistent or aka parallel. just change the y intercept to any number

OpenStudy (anonymous):

so i change the y intercept to forexample 3?

OpenStudy (anonymous):

so i get y=9x+3

OpenStudy (dtan5457):

yes

OpenStudy (dtan5457):

and i think that answers all your questions...unless you have more?

OpenStudy (anonymous):

lol you dont mind if i ask u like what 4 more questions??

OpenStudy (dtan5457):

you'll have to post a new question for each one as stated in the CoC, but tag me and I'll try to get to them.

OpenStudy (anonymous):

ok